OpenStudy (anonymous):

john is x years old and his sister is 24 years younger than him,and his brother is 16 years older than him,if their total age is 100 years,"find each ones age"?

OpenStudy (king):

tell me hw much u hav solved this and where u are stuck??

OpenStudy (anonymous):

im not stuck,i just wanted to see your answers whether it matches mine1

OpenStudy (king):

what is ure answer?

OpenStudy (anonymous):

wait for some time.....maybe about 7 mins!

OpenStudy (king):

why?according to ure previous reply u hav already solved it or r u solving it nw?

OpenStudy (anonymous):

ok....everything is fine..now my ans is quaranteed!

OpenStudy (anonymous):

John's age = x years Sister's age = x-24 years brother's age = x+16 years As per the question, sum of the ages = 100 years.. => x + (x-24) + (x+16) = 100 => x+x-24+x+16 = 100 => 3x -24 + 16 = 100 .. Since, you just want to know the answer, John = 36 years.. Sister = 12 years.. Brother = 52 years .. Do our answers match ?

OpenStudy (king):

u shouldnt have done that aditi nw even if he didnt work out he will say answers match and take the answer....

OpenStudy (anonymous):

hmm does not!

OpenStudy (anonymous):

Haha.. I believed him when he said "im not stuck,i just wanted to see your answers whether it matches mine" .. So.. i did the whole thing :P

OpenStudy (anonymous):

K then, if you don't mind.. show us your working here.. we would find out where did you/I missed out!

OpenStudy (anonymous):

(x)+(x-24)+(x+16)=100

OpenStudy (anonymous):

i did for one....not three!!!!!

OpenStudy (anonymous):

Then wht do u get as x ?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

so 36 +16=bros age and 36-24=sis age right?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

hehe ok thanks

OpenStudy (anonymous):

Welcome :)